OpenStudy (anonymous):

The alternative form of the law of cosines is used to solve for the angle measures of a triangle when given __________. A. All three of the side lengths and no angle measures B. Two of the side lengths and a non-included angle C. Two of the angle measures and no side lengths D. One side length and one angle measure

OpenStudy (anonymous):

Recall that the Law of Cosines is: A^2 = B^2 + C^2 - 2BCcos(α) where α = angle opposite to the side of length A A = length of one side B, C = lengths of other sides To determine the angle measure in the triangle, we need angle to be unknown. Thus, we need three side lengths to be known, which corresponds to choice A.

OpenStudy (anonymous):

okay can you break it down a little more

OpenStudy (anonymous):

Thats kinda as far as I, personally, can go down.

OpenStudy (anonymous):

@Elsa213 @abb0t Can someone else break it down for her?

OpenStudy (anonymous):

okay so the answer is what

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry had a PM message

OpenStudy (anonymous):

np

Elsa213 (elsa213):

>.< Sorry I came late. ;-;